A bus has to cover 72 Km in 1hr 30 minutes. However for the first 36 Km, speed was 54 Km/hr. At what speed must it travel for the rest of distance in order to complete the journey in time.

Answer by mananth(16946)

72 km ------------ 3/2 hour
First part -----36 km----------- 54 km/hr
time = 36/54 = 2/3 hours
Balance time to reach on time = 3/2-2/3= 5/6 hours
Balance distance = 72-36 = 36 km
Speed = distance/time
speed = 36/(5/6)
speed = 36* 6/5
speed = 43.2 km/h second part